OpenStudy (anonymous):

Last year, Xavier purchased 125 shares of Stock A at $96 per share, 30 shares of Stock B at $150 per share, and a $1000 bond with a 4.25% coupon for $1075. Xavier sold both stocks today. Stock A is worth $98 per share and Stock B has a value of $155 per share. Assuming neither stock paid a dividend, which investment has the lowest rate of return?

OpenStudy (anonymous):

Percentage return on stock A is \[\frac{2}{ 96} = 2.1 \%\] Percentage return on stock B is \[\frac{5}{150} = 3.3 \%\] Percentage return on bond (assuming he had not sold it) = ( coupon paid )/1075 = \[\frac{42.5}{1075} =3.95 \%\] so the bond has the highest yield

OpenStudy (anonymous):

thank you!!